A sample of lime CaO weighing 69g prepared by heating 131 g of limestone CaCO3 Calculate the yield of reaction


1
Expert's answer
2021-12-23T05:10:01-0500

CaCO"_3" -> CaO + CO"_2"


Theoretical yield = 131 g "CaCO_3 \times \frac{1 \space mol\space CaCO_3}{100.09g} \times \frac{1 \space mol\space CaO}{1 \space mol CaCo_3} \times \frac{56.08\space g}{1 \space mol}"

Theoretical yield = 73.4 g

% yield = "\frac{69 g }{73.4 g}\times100"

% yield = 94%